OpenStudy (anonymous):

I made this equation, 30x+y=60, is it linear because I need to be linear.

OpenStudy (acxbox22):

a linear equation is where the greatest power in the equation is 1

OpenStudy (acxbox22):

do the variables have a power greater than 1 or less than 1?

OpenStudy (misty1212):

if you want to make it look like \[y=mx+b\] subtract \(30x\) from both sides

OpenStudy (anonymous):

I don't want it to be in slope intercept form

OpenStudy (mrnood):

as long as x AND y do not have powers x^2 or sqrt(y) for eample , i.e. they are both to power 1 then the equation is linear

OpenStudy (freckles):

here are examples of linear equations: 3x+1=5y x=2 4=y 34x+4y+z=1 x+y=1 30x+y=60

OpenStudy (acxbox22):

here is an example of a nonlinear equation x^2+y=10 x^5+x^2+x=234

OpenStudy (acxbox22):

so is it linear?

OpenStudy (mrnood):

it doesn't matter what 'form' it is in - all forms are simply different expressions of the SAME equation. It is linear if the poer of x and y is 1

OpenStudy (anonymous):

I'm kinda confused when an equation is linear or not. What makes an equation linear?

OpenStudy (acxbox22):

the power the variable is raised to @KayleeNeedsHelp

OpenStudy (mrnood):

READ my last post - there is no confusion

OpenStudy (acxbox22):

the power the variable is raised to has to be 1 for it to be linear

OpenStudy (freckles):

\[x^1=x\] usually people don't write the 1 but something like x+y=1 is linear because both x and y are to the first power

OpenStudy (mrnood):

if your equation has no powers (squares, cubes etc. or roots squareroot cuberoot etc.) then it is linear

OpenStudy (anonymous):

@iambatman Could you explain, when I get reply after reply it gets hard to understand anything and @MrNood, I think something has to make a certain type of eqaution, I dont think it would be the same.

OpenStudy (mrnood):

Several people have explained. IF your equation has no powers (squares, cubes etc.) or roots (squareroot cuberoot etc.) then it is linear

OpenStudy (anonymous):

oh well, i guess, Thanks everyone though

OpenStudy (anonymous):

Well thanks to everyone, ill go and carefully read through them. Truely Thanks!! <3

OpenStudy (freckles):

if it is a linear equation with two variables it can written in this form ax+by=c . we can also have a one variable linear equation, x=m or y=n. There are other linear equations with more than two variables. In general it can be written as \[a_nx_n+a_{n-1}x_{n-1}+ \cdots + a_3x_3+a_2x_2+a_1x_1+a_0=d\]

OpenStudy (freckles):

\[m,n,a,b,c,d,a_n ,a_{n-1},...,a_3,a_2,a_1,a_0 \text{ are all constant real values }\]